### v1

@@ 3.3 Inference
Consequently, $N^{sup}$ can be obtained via $\left|\mathcal{I}\right|$. Applying $\mathcal{I}$ to both $\boldsymbol{s}$ and the concatenated $\boldsymbol{v}=\left[\boldsymbol{v}_{1}^{T},\dots,\boldsymbol{v}_{M}^{T}\right]^{T}\in\left\{0,1\right\}^{M\times{N}}$ results in $\boldsymbol{s}^{sup}=\boldsymbol{s}\left[\mathcal{I}\right]\in\left[0,1\right]^{N^{sup}}$ and $\boldsymbol{v}^{sup}=\boldsymbol{v}\left[:,\mathcal{I}\right]\in\left\{0,1\right\}^{M\times{N^{sup}}}$. The vector slicing operator is denoted as $\left[\cdot\right]$. The coarse-grained model score $\boldsymbol{g}\in\mathbb{R}_{+}^{M}$ is computed via $\boldsymbol{v}^{sup}\boldsymbol{s}^{sup}$. The aggregator $\boldsymbol{A}_{agg}=\boldsymbol{A}_{k}$ is selected, where $g_{k}=\left[\boldsymbol{f}_{2}\left(1\right)\right]\left(\boldsymbol{g}\right)$. The support set selected provides a broader knowledge base than the subsequent filtered set. The aggregator is anchored in this larger set to ensure a better understanding of divergent domains, thereby enabling more skilled aggregation.
@@
The hyperparameters $\beta,\,\epsilon,\,\sigma,\,\delta$ are predefined. The filtered indices account for question similarity, model response similarity, and reasoning cost similarity, leading to a more precise support set for better aggregatee selection. With respect to $\mathcal{I}^{flt}$, a fine-grained model score is computed as $\boldsymbol{g}^{f}=\boldsymbol{v}^{sup}\left[:,\mathcal{I}^{flt}\right]\boldsymbol{s}^{flt}\left[\mathcal{I}^{flt}\right]\in\mathbb{R}_{+}^{N}$. Compared with the query-only support set, the filtered set accounts for both queries and LLM responses, helping to filter queries that are semantically similar but differ in difficulty, which is captured by the deep semantics and length of LLM responses.
@@
The $\boldsymbol{f}_{3}:2^{\mathcal{S}}\to\mathcal{S}$ concates any number of strings to a single string. When the number of aggregatees equals 1, aggregation falls back to top-1 fine-grained model score routing. Otherwise, the aggregation dynamically selects LLMs (at most K) to provide context. Note that, to obtain $\bar{x}$, the responses $\hat{y}_{k},\,k\in\left[K\right]$ can be mostly reused due to the overlapping between $\mathcal{A}^{cm}$ and $\mathcal{A}^{fm}$. The adaptive switch fully leverages the advantages of routing and aggregation by filtering noisy LLM responses using a threshold, thereby guaranteeing high-quality aggregation when triggered; otherwise, it provides a more efficient direct-routing path that is less distracting.
@@
DeepSeek-R1-0528 (deepseekai2025deepseekr1incentivizingreasoningcapability) 72.22 64.89 78.33 16.67 76.03 72.97 84.67 28.66 25.33 57.75
DeepSeek-V3-0324 (deepseekai2024deepseekv3technicalreport) 38.89 59.56 68.33 3.70 61.51 59.46 78.44 26.43 24.00 46.70
DeepSeek-V3.1-Terminus (deepseekai2024deepseekv3technicalreport) 55.56 64.67 78.33 8.64 64.67 67.57 84.56 25.12 26.00 52.79
GLM-4.6 (glm-4.6) 88.89 69.56 80.00 14.20 58.99 64.86 80.89 25.89 22.67 56.22
Intern-S1 (bai2025interns1scientificmultimodalfoundation) 38.89 68.00 70.00 9.72 46.69 59.46 83.00 14.33 8.00 44.23
Kimi-K2-0905 (kimiteam2025kimik2openagentic) 72.22 72.22 71.67 5.09 62.15 75.68 80.78 30.66 24.00 54.94
DeepSeek-V3.2-Thinking (deepseekai2025deepseekv32pushingfrontieropen) 88.89 62.44 88.33 24.69 83.91 78.38 87.33 27.81 24.67 62.94
DeepSeek-V3.2-Speciale (deepseekai2025deepseekv32pushingfrontieropen) 94.44 55.33 83.33 27.16 86.75 75.68 87.44 39.52 40.67 65.59
Qwen3-235B-A22B-2507 (yang2025qwen3technicalreport) 77.78 75.33 55.00 9.41 58.36 72.97 83.78 54.01 16.67 55.92
Qwen3-235B-A22B-Thinking-2507 (yang2025qwen3technicalreport) 72.22 77.78 80.00 7.56 75.71 48.65 80.56 49.31 20.00 56.87
@@
Claude-Sonnet-4 (claude-sonnet-4) 41.11 55.47 71.33 4.60 56.85 62.16 83.58 15.58 35.33 47.33
Claude-Sonnet-4.5 (claude-sonnet-4.5) 27.78 64.00 71.67 7.56 60.57 59.46 86.33 16.18 34.00 47.51
Gemini-2.5-Pro (comanici2025gemini25pushingfrontier) 82.22 77.64 84.00 20.62 77.54 45.95 86.69 54.33 34.53 62.61
Grok-4 (grok-4) 88.89 56.89 88.33 24.42 81.03 75.68 86.56 48.38 27.33 64.17
GPT-5 (gpt-5) 83.33 67.11 88.33 25.77 84.54 78.38 87.22 48.00 16.00 64.30
GPT-5.2-Thinking (gpt-5.2) (High Effort) 83.33 85.78 93.33 29.94 90.50 78.38 86.67 35.21 12.67 66.20
Gemini-3-Pro (gemini-3-pro) 94.44 74.55 91.67 33.02 89.59 78.38 89.33 70.03 18.00 71.00
@@
RouterDC (chen2024routerdc) 88.89 76.00 73.33 21.30 83.60 78.38 84.33 53.31 24.00 64.79
EmbedLLM (zhuang2024embedllmlearningcompactrepresentations) 94.44 57.78 90.00 27.31 84.86 75.68 87.44 53.70 42.67 68.21
GraphRouter (feng2025graphroutergraphbasedrouterllm) (Performance-only) 88.89 82.67 90.00 25.93 86.12 78.38 87.33 49.08 27.33 68.41
Avengers (zhang2025avengerssimplerecipeuniting) 94.44 70.44 83.33 27.16 85.49 75.68 87.33 54.08 40.67 68.74
Retrieval-based Router (tang2025opensourcellmscollaborationbeats) 94.44 72.67 86.67 26.85 85.17 75.68 86.78 53.78 40.00 69.12
@@
Multi-agent Methods
@@
MOA (wang2024mixtureofagentsenhanceslargelanguage) (3 Best LLMs) 94.44 80.00 78.33 28.70 88.33 75.68 86.44 42.99 33.33 67.58
Self Consistency (chen2023universalselfconsistencylargelanguage) (Best LLM with 4 Rollouts) 94.44 83.70 85.00 27.62 88.96 75.68 87.56 54.31 38.67 70.66
Majority Voting (chen2024llmcallsneedscaling) (10 LLMs) 83.33 74.67 85.00 13.58 72.56 72.97 87.44 36.98 28.67 61.69
Majority Voting (chen2024llmcallsneedscaling) (3 Best LLMs) 88.89 72.44 86.67 25.93 85.80 75.68 86.78 30.20 29.33 64.64
Our JiSi (w/o Adaptive Aggregation) 94.44 86.44 85.00 30.09 89.27 78.38 87.44 51.46 37.33 71.09
Our JiSi 94.44 88.44 86.67 27.62 89.27 81.08 86.78 53.7 41.33 72.15
Table 1: Main results over diverse benchmarks, LLMs, and methods. The proposed JiSi exhibits superior performance, ranking the best among all open-source and closed-source LLMs, routing methods, and multi-agent methods. $\text{SWE-bench}^{\dagger}$: following the evaluation in (openrouterbench), we pack all context in a single query and drive the LLM or system to solve the issue only via a single response
@@ 4 Experiment
In this section, a comprehensive evaluation of the proposed framework is presented, involving nine benchmarks, fifteen LLMs, five router baselines, and five multi-agent baselines.
@@ 4.1 Settings
Data Preparation. To verify the effectiveness of the proposed framework across benchmarks and different LLM compositions, experiments rely heavily on OpenRouterBench (openrouterbench) (until commit b42d10c) to form both the embedding bank and the test set. For each dataset, the data are split into a training set and a test set in a 0.7:0.3 ratio. In more detail, all data are first added to a list in ascending order by the recorded index, and then shuffled using Numpy (harris2020array) with seed 42. The training set comprises the first 0.7 proportion of the shuffled list, and the test set comprises the remainder. The following benchmarks are involved in the experiments: AIME (aime), Arena-Hard (arenahard2024), GPQA (rein2024gpqa), HLE (phan2025humanitysexam), LiveCodeBench (jain2024livecodebench), LiveMathBench (livemathbench), MMLU-Pro (wang2024mmlu), SimpleQA (wei2024measuringshortformfactualitylarge), and SWE-bench (jimenez2024swebench).
Evaluation. The evaluation follows (openrouterbench) and uses non-interactive mode with a single-round response. Specifically, the judger for Arena-Hard is DeekSeek-V3-0324 (deepseekai2024deepseekv3technicalreport), and the judger for HLE and SimpleQA is o3-mini (o3-mini). For SWE-bench, because aggregation methods are hard to conduct multi-turn interaction and tool usage, we pack all context in a single query and drive the LLM/system to solve the issue only via a single response following (openrouterbench). For an LLM, its responses on the SWE-bench test set are packaged and sent to sb-cli to finalize the results.
@@
Implementation Details. For each item in the training set, the original question and LLM responses are extracted and fed into gte-Qwen2-7B-instruct (li2023generaltextembeddingsmultistage) to obtain embeddings. Additionally, information about completion tokens and response correctness is used to compute reasoning costs and the capability vector. If a response is missing during embedding bank construction or test set inference, an API remote call is triggered to retrieve it from LLM providers, given temperature 0.2 and top-p 1.0. The received response is then cached back to OpenRouterBench. The router and multi-agent methods are applied to ten open-source LLMs ($M=10$), including DeepSeek-R1-0528 (deepseekai2025deepseekr1incentivizingreasoningcapability), DeepSeek-V3-0324 (deepseekai2024deepseekv3technicalreport), DeepSeek-V3.1-Terminus (deepseekai2024deepseekv3technicalreport), GLM-4.6 (glm-4.6), Intern-S1 (bai2025interns1scientificmultimodalfoundation), Kimi-K2-0905 (kimiteam2025kimik2openagentic), DeepSeek-V3.2-Thinking (deepseekai2025deepseekv32pushingfrontieropen), DeepSeek-V3.2-Speciale (deepseekai2025deepseekv32pushingfrontieropen), Qwen3-235B-A22B-2507 (yang2025qwen3technicalreport), and Qwen3-235B-A22B-Thinking-2507 (yang2025qwen3technicalreport). Besides, seven closed-source LLMs are selected for comparison: Claude-Sonnet-4 (claude-sonnet-4), Claude-Sonnet-4.5 (claude-sonnet-4.5), Grok-4 (grok-4), Gemini-2.5-Pro (comanici2025gemini25pushingfrontier), GPT-5 (gpt-5), GPT-5.2-Thinking (gpt-5.2) with high thinking effort, and Gemini-3-Pro (gemini-3-pro). Without otherwise mentioned, $K,\,N^{base},\,\beta,\,\gamma,\,\epsilon,\,\sigma,\,\delta,\,t$ are set to 3, 50, 0.5, 0.95, 0.5, 0.3, 0.2, 0.8, respectively. To avoid extremely long context, the aggregatee number is truncated to 2 if the token sum of 3 aggregatees exceeds 13,000. For computing the cost, if the official API is available, we use the official API price. Otherwise, we adopt the stable API servers with the lowest price in OpenRouter [^2] https://openrouter.ai/.

### v2

@@ 3.3 Inference
Given a new query, the inference phase is composed of three parts: 1) Support-Set-based Aggregator Selection: retrieve $N^{sup}$ queries as the support set based on query embedding similarity; compute coarse-grained model score on the support set; select the aggregator with top-1 coarse-grained model score. 2) Query-Response Mixed Routing: select candidate models with top-K coarse-grained model scores; obtain responses from candidate models with reasoning costs; filter support set to size $N^{flt}$ w.r.t. question, response, and reasoning-cost similarity; compute fine-grained model score on the filtered support set; 3) Adaptive Routing-Aggregation Switch: conduct adaptive aggregation based on fine-grained model score to obtain the final response. The following sections will introduce this procedure in depth.
Given a new query, the inference phase is composed of three parts: 1) Support-Set-based Aggregator Selection; 2) Query-Response Mixed Routing; 3) Adaptive Routing-Aggregation Switch. The following sections will introduce this procedure in depth.
@@
Consequently, $N^{sup}$ can be obtained via $\left|\mathcal{I}\right|$. Applying $\mathcal{I}$ to both $\boldsymbol{s}$ and the concatenated $\boldsymbol{v}=\left[\boldsymbol{v}_{1}^{T},\dots,\boldsymbol{v}_{M}^{T}\right]^{T}\in\left\{0,1\right\}^{M\times{N}}$ results in $\boldsymbol{s}^{sup}=\boldsymbol{s}\left[\mathcal{I}\right]\in\left[0,1\right]^{N^{sup}}$ and $\boldsymbol{v}^{sup}=\boldsymbol{v}\left[:,\mathcal{I}\right]\in\left\{0,1\right\}^{M\times{N^{sup}}}$. The vector slicing operator is denoted as $\left[\cdot\right]$. The coarse-grained model prior score $\boldsymbol{g}\in\mathbb{R}_{+}^{M}$ is computed via $\boldsymbol{v}^{sup}\boldsymbol{s}^{sup}$. The aggregator $\boldsymbol{A}_{agg}=\boldsymbol{A}_{k}$ is selected, where $g_{k}=\left[\boldsymbol{f}_{2}\left(1\right)\right]\left(\boldsymbol{g}\right)$. The support set selected provides a broader knowledge base than the subsequent filtered set. The aggregator is anchored in this larger set to ensure a better understanding of divergent domains, thereby enabling more skilled aggregation.
@@
The hyperparameters $\beta,\,\epsilon,\,\sigma,\,\delta$ are predefined. The filtered indices account for question similarity, model response similarity, and reasoning cost similarity, leading to a more precise support set for better aggregatee selection. With respect to $\mathcal{I}^{flt}$, a fine-grained model prior score is computed as $\boldsymbol{g}^{f}=\boldsymbol{v}^{sup}\left[:,\mathcal{I}^{flt}\right]\boldsymbol{s}^{flt}\left[\mathcal{I}^{flt}\right]\in\mathbb{R}_{+}^{M}$. Compared with the query-only support set, the filtered set accounts for both queries and LLM responses, helping to filter queries that are semantically similar but differ in difficulty, which is captured by the deep semantics and length of LLM responses.
@@
The $\boldsymbol{f}_{3}:2^{\mathcal{S}}\to\mathcal{S}$ concates any number of strings to a single string. When the number of aggregatees equals 1, aggregation falls back to top-1 fine-grained model prior score routing. Otherwise, the aggregation dynamically selects LLMs (at most K) to provide context. Note that, to obtain $\bar{x}$, the responses $\hat{y}_{k},\,k\in\left[K\right]$ can be mostly reused due to the overlapping between $\mathcal{A}^{cm}$ and $\mathcal{A}^{fm}$. The adaptive switch fully leverages the advantages of routing and aggregation by filtering noisy LLM responses using a threshold, thereby guaranteeing high-quality aggregation when triggered; otherwise, it provides a more efficient direct-routing path that is less distracting.
Table 1: Main results over diverse benchmarks, LLMs, and methods. The proposed JiSi exhibits superior performance, ranking the best among all open-source and closed-source LLMs, routing methods, and multi-agent methods. $\text{SWE-bench}^{\dagger}$: following the evaluation in (Li et al., 2026), we pack all context in a single query and drive the LLM or system to solve the issue only via a single response
@@
DeepSeek-R1-0528 (DeepSeek-AI et al., 2025a) 72.22 64.89 78.33 16.67 76.03 72.97 84.67 28.66 25.33 57.75
DeepSeek-V3-0324 (DeepSeek-AI, 2024) 38.89 59.56 68.33 3.70 61.51 59.46 78.44 26.43 24.00 46.70
DeepSeek-V3.1-Terminus (DeepSeek-AI, 2024) 55.56 64.67 78.33 8.64 64.67 67.57 84.56 25.12 26.00 52.79
GLM-4.6 (Z.ai, 2025) 88.89 69.56 80.00 14.20 58.99 64.86 80.89 25.89 22.67 56.22
Intern-S1 (Bai et al., 2025) 38.89 68.00 70.00 9.72 46.69 59.46 83.00 14.33 8.00 44.23
Kimi-K2-0905 (Team et al., 2025) 72.22 72.22 71.67 5.09 62.15 75.68 80.78 30.66 24.00 54.94
DeepSeek-V3.2-Thinking (DeepSeek-AI et al., 2025b) 88.89 62.44 88.33 24.69 83.91 78.38 87.33 27.81 24.67 62.94
DeepSeek-V3.2-Speciale (DeepSeek-AI et al., 2025b) 94.44 55.33 83.33 27.16 86.75 75.68 87.44 39.52 40.67 65.59
Qwen3-235B-A22B-2507 (Yang et al., 2025) 77.78 75.33 55.00 9.41 58.36 72.97 83.78 54.01 16.67 55.92
Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025) 72.22 77.78 80.00 7.56 75.71 48.65 80.56 49.31 20.00 56.87
@@
Claude-Sonnet-4 (Anthropic, 2025a) 41.11 55.47 71.33 4.60 56.85 62.16 83.58 15.58 35.33 47.33
Claude-Sonnet-4.5 (Anthropic, 2025b) 27.78 64.00 71.67 7.56 60.57 59.46 86.33 16.18 34.00 47.51
Gemini-2.5-Pro (Comanici et al., 2025) 82.22 77.64 84.00 20.62 77.54 45.95 86.69 54.33 34.53 62.61
Grok-4 (xAI, 2025) 88.89 56.89 88.33 24.42 81.03 75.68 86.56 48.38 27.33 64.17
GPT-5 (OpenAI, 2025b) 83.33 67.11 88.33 25.77 84.54 78.38 87.22 48.00 16.00 64.30
GPT-5.2-Thinking (OpenAI, 2025a) (High Effort) 83.33 85.78 93.33 29.94 90.50 78.38 86.67 35.21 12.67 66.20
Gemini-3-Pro (Google, 2025) 94.44 74.55 91.67 33.02 89.59 78.38 89.33 70.03 18.00 71.00
@@
RouterDC (Chen et al., 2024b) 88.89 76.00 73.33 21.30 83.60 78.38 84.33 53.31 24.00 64.79
EmbedLLM (Zhuang et al., 2024) 94.44 57.78 90.00 27.31 84.86 75.68 87.44 53.70 42.67 68.21
GraphRouter (Feng et al., 2025) (Performance-only) 88.89 82.67 90.00 25.93 86.12 78.38 87.33 49.08 27.33 68.41
Avengers (Zhang et al., 2025b) 94.44 70.44 83.33 27.16 85.49 75.68 87.33 54.08 40.67 68.74
Retrieval-based Router (Tang et al., 2025) 94.44 72.67 86.67 26.85 85.17 75.68 86.78 53.78 40.00 69.12
@@
Multi-LLM/Multi-Response Methods
@@
MOA (Wang et al., 2024a) (3 Best LLMs) 94.44 80.00 78.33 28.70 88.33 75.68 86.44 42.99 33.33 67.58
Self Consistency (Chen et al., 2023) (Best LLM with 4 Rollouts) 94.44 83.70 85.00 27.62 88.96 75.68 87.56 54.31 38.67 70.66
Majority Voting (Chen et al., 2024a) (10 LLMs) 83.33 74.67 85.00 13.58 72.56 72.97 87.44 36.98 28.67 61.69
Majority Voting (Chen et al., 2024a) (3 Best LLMs) 88.89 72.44 86.67 25.93 85.80 75.68 86.78 30.20 29.33 64.64
Symbolic-MoE (Chen et al., 2025b) 94.44 73.11 83.33 27.47 88.96 75.68 87.00 52.70 19.33 66.89
Self-MoA (Li et al., 2025b) 88.89 80.44 86.67 30.09 84.54 75.68 87.00 55.39 10.00 66.52
JiSi (Ours) 94.44 88.44 86.67 27.62 89.27 81.08 86.78 53.70 41.33 72.15
@@ 4.1 Settings
Data Preparation. To verify the effectiveness of the proposed framework across benchmarks and different LLM compositions, experiments rely heavily on LLMRouterBench (Li et al., 2026) (until commit b42d10c) to form both the embedding bank and the test set. For each dataset, the data are split into a training set and a test set in a 0.7:0.3 ratio. Notably, JiSi is training-free, and the training set is used to construct the embedding bank. The following benchmarks are involved in the experiments: AIME (MAA, 2024), Arena-Hard (Li et al., 2024b), GPQA (Rein et al., 2024), HLE (Phan and others, 2025), LiveCodeBench (Jain et al., 2024), LiveMathBench (Liu et al., 2024), MMLU-Pro (Wang et al., 2024b), SimpleQA (Wei et al., 2024), and SWE-bench (Jimenez et al., 2024).
@@
Implementation Details. For each training query, the original question and LLM responses are extracted and fed into gte-Qwen2-7B-instruct (Li et al., 2023) to obtain embeddings. Information about completion tokens and response correctness is used to compute reasoning costs and the capability vector. The router and multi-LLM methods are applied to ten open-source LLMs ($M=10$), including DeepSeek-R1-0528 (DeepSeek-AI et al., 2025a), DeepSeek-V3-0324 (DeepSeek-AI, 2024), DeepSeek-V3.1-Terminus (DeepSeek-AI, 2024), GLM-4.6 (Z.ai, 2025), Intern-S1 (Bai et al., 2025), Kimi-K2-0905 (Team et al., 2025), DeepSeek-V3.2-Thinking (DeepSeek-AI et al., 2025b), DeepSeek-V3.2-Speciale (DeepSeek-AI et al., 2025b), Qwen3-235B-A22B-2507 (Yang et al., 2025), and Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025). Seven closed-source LLMs are selected for comparison: Claude-Sonnet-4 (Anthropic, 2025a), Claude-Sonnet-4.5 (Anthropic, 2025b), Grok-4 (xAI, 2025), Gemini-2.5-Pro (Comanici et al., 2025), GPT-5 (OpenAI, 2025b), GPT-5.2-Thinking (OpenAI, 2025a) with high thinking, and Gemini-3-Pro (Google, 2025). See Appendix D for details.
4.2 Main Results
As demonstrated in Table 1, the proposed JiSi is compared with the aforementioned closed-source and open-source LLMs, along with typical router methods and multi-LLM methods, i.e., RouterDC (Chen et al., 2024b), EmbedLLM (Zhuang et al., 2024), GraphRouter (Feng et al., 2025), Avengers (Zhang et al., 2025b), Retrieval-based Router (Tang et al., 2025), MoA (Wang et al., 2024a), Self Consistency (Chen et al., 2023), Symbolic-MoE (Chen et al., 2025b), Self-MoA (Li et al., 2025b) and Majority Voting (Chen et al., 2024a). For multi-LLM baselines, priors from LLMRouterBench are applied to provide a stronger variant. For each benchmark, the MoA selects three LLMs with the highest evaluation scores over the training records on the benchmark; the Self Consistency sample four rollouts from the best LLM evaluated on training data and merges responses via majority voting; the Majority Voting is verified in two settings. The first is vanilla, encompassing all open-source LLMs, whereas the second comprises only the top-3 LLMs based on the training set. Besides, the best result of open-source LLMs on the test set of each benchmark is gathered to form the model entry “Best LLM” in Table 1.
Superiority of JiSi. The proposed JiSi framework exhibits promising performance, ranking first on average among all closed-source LLMs, open-source LLMs, router methods, and multi-LLM methods. The performance gain over the most powerful open-source LLM (DeepSeek-V3.2-Speciale) is substantially 6.56%. Even compared with the state-of-the-art closed-source LLM Gemini-3-Pro, a +1.15% gain is achieved. Moreover, the router-only variant of JiSi, which directly adopts $\boldsymbol{A}_{1}^{fm}$ without aggregation formulated in Eq. 5, outperforms other router methods, indicating that the proposed query-response mixed router can yield more suitable aggregatees for subsequent aggregation. Compared with the router-only variant, JiSi surpasses the best method by +3.03%. Besides, JiSi overpasses the strong performance bound of routing, i.e. “Best LLM”, by +1.6%, revealing enormous potential. Meanwhile, we also conduct the ablation study of JiSi in Appendix C.
Efficiency of JiSi. As shown in Table 2, we report the cost of the proposed JiSi and the leading proprietary LLMs. It is worth noting that although JiSi utilizes more API calls due to multi-LLM aggregation, it can remarkably reduce the total cost while maintaining superior performance compared with leading proprietary LLMs. For instance, JiSi surpasses Gemini-3-Pro and Grok-4 by +1.15% and +7.89%, while saving 53.23% and 81.68% cost. The reasons lie in two aspects: 1) the APIs of open-source LLMs are dramatically cheaper than proprietary LLMs. 2) The adaptive routing-aggregation switch mitigates the workload of aggregation and even simplifies aggregation to routing. The strong trade-off between efficiency and effectiveness demonstrates the feasibility and economy of JiSi in practical implementation. See Appendix D.2 for the detailed LLM prices.